OpenStudy (anonymous):

a solid rectangular brick is to be made from 1 cu feet of clay. the brick must be 3 times as long as wide the width of brick for which it will have minimum surface area is a find a^3

OpenStudy (anonymous):

if the dimensions of rectangular brick are x,a,y then x*a*y=1 and x=3a and what else can be done are will be 2(xy+ya+ax) and to minimise it i need to differentiate it and put it zero i did it but the answer wont match so pl could anyone help me

ganeshie8 (ganeshie8):

wat did u get for a^3 ?

OpenStudy (anonymous):

i got something like this 2(18a^3+3a^2-1)=0 i cant simplify it further the options given are (2/9),(2/9)^(1/3),8/729,3/2

ganeshie8 (ganeshie8):

im getting 2/9

ganeshie8 (ganeshie8):

surface area : 2(xy+ya+ax) substitute x and y values in terms of a

ganeshie8 (ganeshie8):

x = 3a ***** xay = 1 3a^2y = 1 y = 1/(3a^2) ****

ganeshie8 (ganeshie8):

surface area : 2(1/a + 1/(3a) + 3a^2) minimize this

OpenStudy (anonymous):

ok thanks i was doing it wrong i got the surface area wrong

ganeshie8 (ganeshie8):

happens... np :)

OpenStudy (anonymous):

hey but how did u solved the cubic that we got on differentiating

ganeshie8 (ganeshie8):

u wil get an easy expression, pen down and see...

OpenStudy (anonymous):

s=2(1/a+1/(3a)+3a^2) s'=0=2(-1/a^2 -1/(3a^2) +6a) -1-1/3+6a^3=0 so 4/3=6a^3 so a^3=2/9 itsnt it?

ganeshie8 (ganeshie8):

looks good !

OpenStudy (anonymous):

thanks a lot

ganeshie8 (ganeshie8):

u wlc :)